Find the distance and time at which a fast moving train B will excell the train A. Where they are moving in the same direction with velocity 50 km/hr for A and 80 km/hr for B. Where train A has initial displacement of 1200 km & train B has no initial displacement
1
Expert's answer
2020-10-15T03:19:03-0400

solution

velocity of train A(va)=50km/hr

velocity of train B(vb)=80km/hr

displacement of train A(sa)=1200 km


velocity of train B with respect to train A(vba) =vb - va


vba =80 - 50=30km/hr


time taken by train B to over take the train A(t)="\frac{s_a}{v_{ba}}=\frac{1200}{30}=40hr"


distance travel by train B in 40hr (sb) ="v_bt"

"s_b=80\times40=3200km"


so times taken by train B to excell train A is 40hr and distance travel by train B is 3200km.